OpenStudy (96mertcank96):

what is the product of 4x/x-1 and x^2-1/3x+3 ?

OpenStudy (anonymous):

\[\frac{4x}{x-1}*\frac{x^2-1}{3x+3} = \frac{4x*(x^2-1)}{(x-1)(3)(x+1)}\] (x-1)*(x+1) = x^2-1

OpenStudy (96mertcank96):

Thank You!

OpenStudy (anonymous):

welcome